The expression - IF(A1 > 12, 6*A1, 2*A1) is used in a spreadsheet. Find the result if A1 is 7 Find the result if A1 is 13.
Answer #1

The expression can be written in layman's terms as follows

If the value of A1 is more than 12, then the value of the result will be 6*A1

If the value of A1 is not more than 12, then the value of the result will be 2*A1

Case 1:
A1=7

The value of A1, as we can see, is not more than 12

Hence, the value of the result will be 2*A1 , that is 2*7=14

Case 2:

A1=13

The value of A1, as we can see, is more than 12

Hence, the value of the result will be  6*A1=6*13=78